# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 5.3. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ and the ring of cooperations
    It was shown in [CQ21, Lemma 3.38] that the groups $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$ are not expressible in terms of Adams covers of kq and ksp. The purpose of the groups $Z_{i}^{\mathbb{C}}$ is exactly to account for this failure, mixing in an appropriate trigraded version of $\text{Ext}_{\mathcal{A}(1)^{\vee}}(\text{H}_{**}(\text{bo}^{\langle i\rangle}))$ and $\text{Ext}_{\mathcal{A}(1)^{\vee}}(\text{H}_{**}(\text{bsp}^{\langle i\rangle})).$ The difference being observed in the computation of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ can be accounted for by noting that this group also cannot be expressed using just $Z_{2k-\alpha(k)}^{\mathbb{C}}$. Rather, we must express these groups using shifts of $Z_{\alpha(k)}^{\mathbb{C}}$ and an appropriate trigraded version of the $h_{0}$-towers found in the classical groups $\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{F}_{2})$; this last term is precisely what is given by $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}))$ in the case that $k\equiv 0\,(2),$ and by $\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1))$ in the case that $k\equiv 1\,(2).$
@@
[…]
@@
Proposition 5.10.
[…]
$E_{2}=\bigoplus_{k\geq 0}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\Sigma^{4k,2k}B_{0}^{\mathbb{C}}(k))\cong\bigoplus_{k\geq 0}\left(\Sigma^{4k,2i}Z^{\mathbb{C}}_{\alpha(k)}\oplus\bigoplus_{j=0}^{4k-8}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\right).$
[…]
Theorem 5.11.
@@
[…]
$\text{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\textup{H}_{**}(\textup{kq}\otimes\textup{kq}))\to\text{Ext}_{\mathcal{A}(1)^{\vee}}^{s,f}(\text{H}_{*}(\text{bo}\otimes\text{bo})),$
[…]
@@ 6. Algebraic preliminaries and comparisons
6.1. The dual motivic Steenrod algebra and $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$
Let $\mathbb{M}_{2}^{\mathbb{R}}=\mathbb{F}_{2}[\rho,\tau]$ denote the mod-2 motivic homology of a point [Voe03a, Corollary 6.10], where $|\tau|=(0,-1)$ and $|\rho|=(-1,-1)$. Recall from 2.7 that the dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{\mathbb{R}}$ takes the form
@@
$E_{2}^{s,f,w}=\bigoplus_{k\geq 0}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{s,f,w}(\Sigma^{4k,2k}B_{0}^{\mathbb{R}}(k)),$
[…]
Theorem 6.1.
[…]
$\textup{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})=\mathbb{F}_{2}[\tau^{4},\rho,h_{0},h_{1},a,b,\tau^{2}h_{0},\tau h_{1},b,\tau^{2}a]/I.$
The degrees of the generators are listed in Table 1 and the ideal of relations $I$ is detailed in Table 2.
@@
Table 1. Generators for $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$
@@
$\rho\cdot\tau^{2}h_{0}$ $(-1,1,-3)$ 2
$\rho^{3}\cdot a$ $(1,3,-1)$ 2
$\tau^{2}h_{0}\cdot h_{1}+\rho(\tau h_{1})^{2}$ $(1,2,-1)$ 2
@@
Table 2. Relations for $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$
The data comprising $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ is too complicated to display in one chart. To present this material more clearly, we give individual charts for each coweight modulo 4, where this 4-periodicity is induced by the generator $\tau^{4}$. Notice that there are no generators in coweight 3 modulo 4, and the relations ensure that there are never any classes in this coweight. Charts are written in $(s,f)$-grading, with motivic weight suppressed. We also suppress infinite $\rho$-towers, which would extend to the left in the charts, for readability. Fig. 6 depicts the coweight 0 piece, Fig. 7 depicts the coweight 1 piece, and Fig. 8 depicts the coweight 2 piece.
A black $\blacksquare$ represents $\mathbb{F}_{2}[\tau^{4}][\rho,h_{1}]$. A black $\bullet$ represents $\mathbb{F}_{2}[\tau^{4}]$. A vertical black line represents multiplication by $h_{0}$. A horizontal black line represents multiplication by $\rho$. A diagonal black line represents multiplication by $h_{1}$. A dashed horizontal line indicates that $\rho$-multiplication hits $\tau^{4}$ times a generator. For example, in Fig. 6 there is a dashed line indicating $\rho\cdot\tau^{2}a=\tau^{4}h_{1}^{3}$.
Figure 6. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}.}(\mathbb{M}_{2}^{\mathbb{R}})$ in $cw\equiv 0\,(4)$
Figure 7. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}.}(\mathbb{M}_{2}^{\mathbb{R}})$ in $cw\equiv 1\,(4)$
Figure 8. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}.}(\mathbb{M}_{2}^{\mathbb{R}})$ in $cw\equiv 2\,(4)$
6.2. Some modules over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$
We now introduce notation for some modules over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ which will appear throughout. The reader is encouraged to consult Appendix A to follow along with these definitions.
@@ 6.2. Some modules over $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$
Define the following $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
  • Let $\mathcal{D}$ denote the entire coweight $cw\equiv 1\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$, called a big diamond (see Fig. 7);
  • Let $\mathcal{S}$ denote the entire coweight $cw\equiv 2\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, called a big staircase (see Fig. 8);
  • Let $P=\mathbb{F}_{2}[\tau^{4}][\rho]$, called a $\rho$-tower, with the obvious $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Fig. 28);
  • Let $H=\mathbb{F}_{2}[\tau^{4}][h_{0}]$, called an $h_{0}$-tower, with the obvious $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Fig. 29);
  • Let $PH=\mathbb{F}_{2}[\tau^{4}][\rho,h_{0}]/(\rho h_{0})$, called a $(\rho,h_{0})$-tower, with the obvious $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Fig. 30);
  • Let $D=\mathbb{F}_{2}[\tau^{4}][\rho,h_{0},h_{1}]/(\rho^{2},h_{0}^{2},h_{1}^{2},\rho h_{0},h_{0}h_{1},\rho h_{1}=h_{0})$, called a diamond, with the obvious $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Fig. 31);
  • Let $S=\mathbb{F}_{2}[\tau^{4}][h_{0},h_{1},t,a]/(h_{1}^{3},t^{2},a^{2},h_{0}h_{1},h_{1}a,ta,h_{0}^{2}t,h_{1}^{2}t$), called a staircase, with the $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure given by the relations (see Fig. 32)
@@
  • Let $T=\mathbb{F}_{2}[\tau^{4}][\rho]/(\rho^{2}),$ called a segment, with the obvious $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Fig. 33);
  • Let $J=\mathbb{F}_{2}[\tau^{4}][\rho,h_{0}]/(\rho^{3},\rho h_{0})$, called a J-tower, with the obvious $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Fig. 34);
  • Let $JD=\mathbb{F}_{2}[\tau^{4}][\rho,h_{0},t]/(\rho^{3},t^{2},\rho^{2}t,\rho h_{0},h_{0}^{2}t)$, called a $JD$-tower, with the $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure given by the relations (see Fig. 35)
@@
We construct now a more intricate family of modules. This module will track the behavior of the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})$ throughout the process of the coming spectral sequences.
@@
In this construction, we will attach generators to extend the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$. We continue until we have attached a generator in Adams filtration 0, at which point we stop. For $n\geq 0$, let $\mathcal{F}_{s,f,w}$ be the $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module concentrated in coweight $cw\equiv s-w\,(4)$, called a big flag, described by the following process:
@@
    There is a copy of the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ (see Fig. 6) with generator $x_{0}$ in degree $(s,f,w)$ called a flag;
@@
    There is a $PH$ generated by $x_{4n+1}$ in degree $(s-4-8n,f-1-4n,w-4-8n)$ subject to the relations
@@
    There is a $P$ generated by $x_{4n+2}$ in degree $(s-6+8n,f-2-4n,w-6-8n)$ subject to to the relations
@@
    There is a $P$ generated by $x_{4n+3}$ in degree $(s-7-8n,f-3-4n,w-7-8n)$ subject to the relations
@@
    There is a $PH$ generated by $x_{4n+4}$ in degree $(s-8(n+1),f-4(n+1),w-8(n+1))$ subject to the relations
@@
In short, the submodule $\mathcal{F}_{s,f,w}$ looks like the coweight $cw\equiv 0\,(4)$ portion of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ with generator shifted from $(0,0,0)$ to $(s,f,w)$, and then a complicated “tail” pattern coming from the generator of the flag which decreases in all degrees until Adams filtration 0 is reached. Importantly, at each step we are attaching a new generator by an $h_{1}$, and since $h_{1}$ has coweight 0 as an element of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, this ensures that each stage of this process is concentrated in the same coweight modulo 4. As an example, we depict the big flag $\mathcal{F}_{12,3,6}$ in Fig. 9.
Figure 9. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $\mathcal{F}_{12,3,6}$.

### v2

@@ 5.2. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$
    It was shown in [17, Lemma 3.38] that the groups $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$ are not expressible in terms of Adams covers of kq and ksp. The purpose of the groups $Z_{i}^{\mathbb{C}}$ is exactly to account for this failure, mixing in an appropriate trigraded version of $\text{Ext}^{**}_{\mathcal{A}(1)_{\text{cl}}^{\vee}}(\text{H}_{*}(\text{bo}^{\langle i\rangle}))$ and $\text{Ext}^{**}_{\mathcal{A}(1)_{\text{cl}}^{\vee}}(\text{H}_{*}(\text{bsp}^{\langle i\rangle})).$ The difference being observed in the computation of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ can be accounted for by noting that this group also cannot be expressed using just $Z_{2k-\alpha(k)}^{\mathbb{C}}$. Rather, we must express these groups using shifts of $Z_{\alpha(k)}^{\mathbb{C}}$ and an appropriate trigraded version of the $h_{0}$-towers found in the classical groups $\text{Ext}^{**}_{\mathcal{A}(1)^{\vee}_{\text{cl}}}(\mathbb{F}_{2})$; this last term is precisely what is given by $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}))$ in the case that $k\equiv 0\,(2),$ and by $\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1))$ in the case that $k\equiv 1\,(2).$
@@
    Apart from this discrepancy, the contents of [17] are unaffected. In particular, we will see that the $\textbf{mASS}^{\mathbb{C}}(\text{kq}\otimes\text{kq})$ collapses on the $\textup{E}_{2}$-page by using the same arguments as employed by Culver–Quigley.
@@
Proposition 5.6 ([17, Proposition 3.41]).
The $\textup{E}_{2}$-page of the $\textup{{mASS}}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq})$ is given, modulo $v_{1}$-torsion, by
$\textup{E}_{2}=\bigoplus_{k\geq 0}\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\Sigma^{4k,2k}B_{0}^{\mathbb{C}}(k))\cong\bigoplus_{k\geq 0}\left(\Sigma^{4k,2i}Z^{\mathbb{C}}_{\alpha(k)}\oplus\bigoplus_{j=0}^{4k-8}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\right).$
Recall from [41, Theorem 2.9] that the ASS(bobo)($\text$$\otimes\text{bo}$) collapses on $\textup{E}_{2}$.
Theorem 5.7 ([17, Cor 3.43]).
@@
Betti realization sends the mASS($\textup{kq}\otimes\textup{kq})$ to the ASS($\text{bo}\otimes\text{bo}$). At the level of $\textup{E}_{2}$-pages, this is a map
$\text{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\textup{H}_{**}(\textup{kq}\otimes\textup{kq}))\to\text{Ext}_{\mathcal{A}(1)_{\text{cl}}^{\vee}}^{s,f}(\text{H}_{*}(\text{bo}\otimes\text{bo})),$
obtained by inverting $\tau$ on the $\textup{E}_{2}$-page of the mASS($\textup{kq}\otimes\textup{kq}$) and setting $\tau=1$. Since the $\textbf{ASS}(\text{bo}\otimes\text{bo})$ collapses on $\textup{E}_{2}$, we must have that there are no motivic differentials where both source and target are $\tau$-torsion free. However, our description of the $\textup{E}_{2}$-page of the $\textbf{mASS}(\textup{kq}\otimes\textup{kq})$ shows that any $\tau$-torsion class must be $h_{1}$-torsion free, hence is some $h_{1}$-tower. For degree reasons, there can be no such differentials between these towers, concluding the proof. ∎
@@ 6. Algebraic preliminaries and comparisons
6.1. Background
Let $\mathbb{M}_{2}^{\mathbb{R}}=\mathbb{F}_{2}[\rho,\tau]$ denote the mod-2 motivic homology of a point [55, Corollary 6.10], where $|\tau|=(0,-1)$ and $|\rho|=(-1,-1)$. Recall from 2.4 that the dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{\mathbb{R}}$ takes the form
@@ 6.1. Background
4.1 asserts that the $\textbf{mASS}^{\mathbb{R}}(\textup{kq}\otimes\textup{kq})$ takes the form
$\textup{E}_{2}^{s,f,w}=\bigoplus_{k\geq 0}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{s,f,w}(\Sigma^{4k,2k}B_{0}^{\mathbb{R}}(k))\implies\pi_{s,w}^{\mathbb{R}}(\text{kq}\otimes\text{kq}),$
where $B_{0}^{\mathbb{R}}(k)$ are the $\mathbb{R}$-motivic integral motivic Brown–Gitler comodules. Therefore, to compute this $\textup{E}_{2}$-page we must compute the trigraded groups $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(k))$ for $k\geq 0.$ Since $B_{0}^{\mathbb{R}}(0)=\mathbb{M}_{2}^{\mathbb{R}}$, the first summand is isomorphic to the $\textup{E}_{2}$-page of the $\textbf{mASS}^{\mathbb{R}}(\textup{kq})$. This was computed by Hill in [26, Theorem 5.6]. We use a presentation adopted by Guillou–Hill–Isaksen–Ravenel [24, Theorem 6.2].
Theorem 6.1 ([26, 24]).
@@
$\textup{E}_{2}^{s,f,w}=\textup{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}},\mathbb{M}_{2}^{\mathbb{R}})=\mathbb{F}_{2}[\tau^{4},\rho,h_{0},h_{1},a,b,\tau^{2}h_{0},\tau h_{1},b,\tau^{2}a]/I\implies\pi_{s,w}^{\mathbb{R}}(\textup{kq}).$
The spectral sequence collapses at $\textup{E}_{2}$. The degrees of the generators are listed in Table 1 and the ideal of relations II is described in Table 2.
@@
Table 1. Generators for $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$.
@@
$\rho\cdot\tau^{2}h_{0}$ (1,1,3)(-1,1,-3) 2
$\rho^{3}\cdot a$ (1,3,1)(1,3,-1) 2
$\tau^{2}h_{0}\cdot h_{1}+\rho(\tau h_{1})^{2}$ (1,2,1)(1,2,-1) 2
@@
Table 2. Relations for $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$.
The data comprising $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ is too complicated to display in one chart. To present this material more clearly, we give individual charts for each coweight modulo 4, where this 4-periodicity is induced by the generator $\tau^{4}$. Notice that there are no generators in coweight 3 modulo 4, and the relations ensure that there are never any classes in this coweight. Charts are written in $(s,f)$-grading, with motivic weight suppressed. We also suppress infinite $\rho$-towers, which would extend to the left in the charts, for readability. Figure 6.1 depicts the coweight 0 piece, Figure 6.2 depicts the coweight 1 piece, and Figure 6.3 depicts the coweight 2 piece.
A black $\blacksquare$ represents $\mathbb{F}_{2}[\rho,\tau^{4}]$. A black $\bullet$ represents $\mathbb{F}_{2}[\tau^{4}]$. A vertical black line represents multiplication by $h_{0}$. A horizontal black line represents multiplication by $\rho$. A diagonal black line represents multiplication by $h_{1}$. A dashed horizontal line indicates that $\rho$-multiplication hits $\tau^{4}$ times a generator. For example, in Figure 6.1 there is a dashed line indicating $\rho\cdot\tau^{2}a=\tau^{4}h_{1}^{3}$.
Figure 6.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}.}(\mathbb{M}_{2}^{\mathbb{R}})$ in $cw\equiv 0\,(4)$.
Figure 6.2. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}.}(\mathbb{M}_{2}^{\mathbb{R}})$ in $cw\equiv 1\,(4)$.
Figure 6.3. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}.}(\mathbb{M}_{2}^{\mathbb{R}})$ in $cw\equiv 2\,(4)$.
6.2. Some modules over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$
We now introduce notation for some modules over $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ which will appear throughout. The reader is encouraged to consult Appendix A to follow along with these definitions.
@@
Define the following $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-modules:
  • Let $\mathcal{D}$ denote the entire coweight $cw\equiv 1\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{R}})$, called a big diamond (see Figure 6.2);
  • Let $\mathcal{S}$ denote the entire coweight $cw\equiv 2\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, called a big staircase (see Figure 6.3);
  • Let $P$ be the module
    $P=\mathbb{F}_{2}[\rho,\tau^{4}],$
    called a $\rho$-tower, with the obvious $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Figure A.4);
  • Let $H$ be the module
    $H=\mathbb{F}_{2}[h_{0},\tau^{4}],$
    called an $h_{0}$-tower, with the obvious $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Figure A.4);
  • Let $PH$ be the module
    $PH=\frac{\mathbb{F}_{2}[h_{0},\rho,\tau^{4}]}{(\rho h_{0})},$
    called a $(\rho,h_{0})$-tower, with the obvious $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Figure A.4);
  • Let $D$ be the module
    $D=\frac{\mathbb{F}_{2}[h_{0},h_{1}\rho,\tau^{4}]}{(\rho^{2},h_{0}^{2},h_{1}^{2},\rho h_{0},h_{0}h_{1},\rho h_{1}=h_{0})},$
    called a diamond, with the obvious $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Figure A.4);
  • Let $S$ be the module
    $S=\frac{\mathbb{F}_{2}[h_{0},h_{1},t,a,\tau^{4}]}{(h_{1}^{3},t^{2},a^{2},h_{0}h_{1},h_{1}a,ta,h_{0}^{2}t,h_{1}^{2}t)},$
    called a staircase, with $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure given by the relations (see Figure A.8)
@@
  • Let $T$ be the module
    $T=\frac{\mathbb{F}_{2}[\rho,\tau^{4}]}{(\rho^{2})},$
    called a segment, with the obvious $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Figure A.8);
  • Let $J$ be the module
    $J=\frac{\mathbb{F}_{2}[h_{0},\rho,\tau^{4}]}{(\rho^{3},\rho h_{0})},$
    called a J-tower, with the obvious $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure (see Figure A.8);
  • Let
    $JD=\frac{\mathbb{F}_{2}[h_{0},t,\rho,\tau^{4}]}{(\rho^{3},t^{2},\rho^{2}t,\rho h_{0},h_{0}^{2}t)},$
    called a JDJD-tower, with $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module structure given by the relations (see Figure A.8)
@@
We construct now a more intricate family of modules. This module will track the behavior of the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ throughout the coming spectral sequences.
@@
In this construction, we will attach generators to extend the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$. We continue until we have attached a generator in Adams filtration 0, at which point we stop. For $n\geq 0$, let $\mathcal{F}_{s,f,w}$ be the $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module concentrated in coweight $cw\equiv s-w\,(4)$, called a big flag, described by the following process:
@@
    There is a copy of the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ (see Figure 6.1) with generator $x_{0}$ in degree $(s,f,w)$ called a flag.
@@
    There is a $PH$ generated by $x_{4n+1}$ in degree (s48n,f14n,w48n)(s-4-8n,f-1-4n,w-4-8n) subject to the relations
@@
    There is a $P$ generated by $x_{4n+2}$ in degree (s68n,f24n,w68n)(s-6-8n,f-2-4n,w-6-8n) subject to to the relations
@@
    There is a $P$ generated by $x_{4n+3}$ in degree (s78n,f34n,w78n)(s-7-8n,f-3-4n,w-7-8n) subject to the relations
@@
    There is a $PH$ generated by $x_{4n+4}$ in degree (s8(n+1),f4(n+1),w8(n+1))(s-8(n+1),f-4(n+1),w-8(n+1)) subject to the relations
@@
In short, the submodule $\mathcal{F}_{s,f,w}$ looks like the coweight $cw\equiv 0\,(4)$ portion of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ with generator shifted from $(0,0,0)$ to $(s,f,w)$, and then a complicated “tail” pattern coming from the generator of the flag which decreases in all degrees until Adams filtration 0 is reached. Importantly, at each step we are attaching a new generator by $h_{1}$, and since $h_{1}$ has coweight 0 as an element of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, this ensures that each stage of this process is concentrated in the same coweight modulo 4. As an example, we depict the big flag $\mathcal{F}_{12,3,6}$ in Figure 6.4.
Figure 6.4. The $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$-module $\mathcal{F}_{12,3,6}$.